OpenStudy (anonymous):

I need help with the steps! solve by elimination; x+10y=-10 8x+5y=-5 please/thanks:)

OpenStudy (nottim):

Rearrange one of them to solve for either x or y, and then sub it into the other equation.

OpenStudy (anonymous):

2times (8x+5y=-5) - (x+10y=-10) <--- get x

OpenStudy (anonymous):

How about you multiply the bottom equation by -2 so that when you add them, your y variables cancel?

OpenStudy (nottim):

So, for example. The top one, rearranged to solve for x is x=-10y-10. Sub that into the bottom equation to get 8(-10y-10)+5y=-5

OpenStudy (anonymous):

Thanks guys, that helped!

OpenStudy (nottim):

No Problem, citizen! UP UP AND A BATMAN AWAT

OpenStudy (nottim):

away

OpenStudy (callisto):

@NotTim i think that is substitution, not elimination

OpenStudy (nottim):

OH BOTHER. YOU'RE RIGHT.

OpenStudy (nottim):

BUGGER. How do I do it by elimination, I shan't remember.

OpenStudy (anonymous):

ahhh crap im confused now

OpenStudy (callisto):

For elimination, multiply a constant on both sides of a equation so that you can cancel it by adding to / subtracting from the other one. Take the question as an example. x+10y=-10 -(1) 8x+5y=-5 -(2) the easier way is to multiply (2) by 2 and subtract it from (1) that is (1) - (2)x2 (x+10y) - (8x+5y)x2 = (-10) -(-5)x2 you will see that you can cancel the unknown y after expanding it

OpenStudy (callisto):

(x+10y) - (16x + 10y) = (-10) + 10 x + 10y -16x -10y = 0 and solve it..

OpenStudy (anonymous):

THANK YOU @Callisto

OpenStudy (callisto):

Do you really understand?

OpenStudy (anonymous):

At least this problem yes!

OpenStudy (callisto):

Then try another way by cancelling x for this question

OpenStudy (anonymous):

blehhhhhhh:(

OpenStudy (callisto):

i need to make sure you really understand the skill...

OpenStudy (anonymous):

:( FINE! but im not gonna like it-___-

OpenStudy (anonymous):

i dont get it:O